# A Bernstein polynomial approach for the estimation of cumulative distribution functions in the presence of missing data

### v1

@@ Abstract
We study nonparametric estimation of univariate cumulative distribution functions (CDFs) pertaining to data missing at random. The proposed estimators smooth the inverse probability weighted (IPW) empirical CDF with the Bernstein operator, yielding monotone, $[0,1]$-valued curves that automatically adapt to bounded supports. We analyze two versions: a pseudo estimator that uses known propensities and a feasible estimator that uses propensities estimated nonparametrically from discrete auxiliary variables, the latter scenario being much more common in practice. For both, we derive pointwise bias and variance expansions, establish the optimal polynomial degree $m$ with respect to the mean integrated squared error, and prove the asymptotic normality. A key finding is that the feasible estimator has a smaller variance than the pseudo estimator by an explicit nonnegative correction term. We also develop an efficient degree selection procedure via least-squares cross-validation. Monte Carlo experiments demonstrate that, for moderate to large sample sizes, the Bernstein-smoothed feasible estimator outperforms both its unsmoothed counterpart and an integrated version of the IPW kernel density estimator proposed by Dubnicka, (2009) in the same context. A real-data application to fasting plasma glucose from the 2017–2018 NHANES survey illustrates the method in a practical setting. All code needed to reproduce our analyses is readily accessible on GitHub.
@@ 1 Introduction
The cumulative distribution function (CDF) and its inverse (the quantile function) are fundamental objects in probability and statistics, providing a complete description of a random variable’s distribution. They offer insights into the entire distribution, including tail behavior, that summary measures like the mean or variance cannot capture. Accurate estimation of CDFs is crucial across numerous fields, for example, in survival analysis (where the CDF relates to survival probabilities), reliability engineering (where the CDF gives failure probabilities and reliability over time), economics (where the CDF characterizes income, wealth, or return distributions), and risk management (where the loss CDF underpins tail-risk metrics such as Value-at-Risk), where one often needs to understand the full distribution rather than just a single parameter.
@@
In practice, statistical analysis is frequently complicated by incomplete data. Missing observations are ubiquitous in large studies, clinical trials, and surveys. If data are missing in a nontrivial fashion, standard nonparametric estimators such as the empirical CDF computed from observed cases can be biased for the true distribution unless the data are missing completely at random, that is, the missingness mechanism is independent of the data values (Little & Rubin,, 2002, p.12). A more plausible assumption in many contexts is missing at random (MAR), which posits that the probability an observation is missing may depend on fully observed auxiliary variables but not on the value of the missing observation itself (Rubin,, 1976). Under MAR, one can obtain unbiased estimates by appropriately reweighting or imputing the observed data. A prominent approach is inverse probability weighting (IPW), inspired by the Horvitz–Thompson estimator from survey sampling (Horvitz & Thompson,, 1952). In IPW, each observed case is weighted by the inverse of its observation probability (called propensity score), often estimated from covariates (Rosenbaum & Rubin,, 1983), yielding a weighted empirical CDF (sometimes called a pseudo-empirical CDF). When propensity scores are known or consistently estimated, this IPW empirical CDF is asymptotically unbiased for the full-population CDF.
@@
Alternative nonparametric strategies for MAR data include imputation-based methods and hybrid approaches. For instance, Cheng & Chu, (1996) proposed a kernel regression imputation to consistently estimate the CDF and quantiles, proving strong uniform consistency and asymptotic normality. Hybrid strategies, in particular augmented IPW estimators, have been developed to improve efficiency by combining weighting with outcome imputation. Wang & Qin, (2010) introduced an augmented IPW empirical-likelihood approach for CDFs with missing responses that delivers asymptotic Gaussian limits for the process. These methods underscore that, under MAR assumptions, one can recover distributional information using either weighting or imputation, or both, without fully specifying outcome models. Furthermore, these estimation strategies have also been adapted to the more complex case of nonignorable missing data (NMAR). Related work by Ding & Tang, (2018) considered IPW, regression imputation, and an augmented approach for distribution and quantile estimation under a semiparametric NMAR model, finding that all three can reach the same asymptotic variance under suitable conditions.
@@
However, the IPW-based empirical CDF, like the ordinary empirical CDF, is a step function. When the underlying distribution is continuous, it is often desirable to smooth such step-function estimates to improve efficiency, reduce mean squared error (MSE), and produce a smooth, continuous CDF estimate. Smoothing can yield substantial gains; some smooth CDF estimators have been shown to asymptotically outperform the raw empirical CDF in mean integrated squared error (MISE). Kernel smoothing is a popular approach in the density estimation context; see, e.g., Rosenblatt, (1956); Parzen, (1962), or Silverman, (1986); Wand & Jones, (1995) for book treatments. This approach extends naturally to CDFs by integrating the kernel density estimate, resulting in the classical smooth CDF estimator introduced by Nadaraya, (1964), which replaces the indicator steps of the empirical CDF with integrated kernels. An alternative strategy involves direct regression smoothing of the empirical CDF. For example, Cheng & Peng, (2002) proposed a local linear estimator that can achieve a smaller asymptotic MISE than the integrated kernel estimator, although it does not guarantee monotonicity. These kernel methods have been adapted to handle missing-data scenarios; for example, Dubnicka, (2009) developed an IPW kernel density estimator (KDE) under MAR. For CDFs under MAR, a natural idea is to smooth the IPW variant using similar techniques. For instance, one of the competitors considered in our Monte Carlo study (Section 4) is an integrated version of Dubnicka’s estimator.
@@
Smoothing the CDF can improve pointwise variance and yield well-behaved estimates, but one must address the well-known boundary bias that traditional kernel estimators exhibit on bounded supports; see, e.g., Zhang et al. , (2020). If the support of the distribution is [0,1], integrated-KDE CDF estimators tend to incur bias near 0 and 1 due to a spill-over effect created by the fixed kernel. To mitigate this, various boundary-correction techniques have been proposed, including reflecting the data at boundaries or using boundary kernels tailored for the edges (Tenreiro,, 2013). A user-friendly and highly effective strategy employs asymmetric kernels, which inherently respect support limits and ensure monotonicity by construction; see, e.g., Lafaye de Micheaux & Ouimet, (2021); Mansouri et al. , (2023). Alternatively, monotonicity constraints can be imposed on smoothing techniques that do not naturally guarantee it. For instance, Xue & Wang, (2010) proposed a constrained polynomial spline approach that smooths the empirical CDF while enforcing monotonicity, thereby guaranteeing a valid CDF estimate.
@@
In this paper, the focus is on CDFs supported on the unit interval $[0,1]$ (more general bounded supports can often be transformed to $[0,1]$). For such cases, Bernstein polynomials offer an elegant and effective smoothing technique. Bernstein polynomials (Bernstein,, 1912) were originally introduced as a constructive proof of the Weierstrass approximation theorem (Weierstraß,, 1885), which asserts the existence of uniform polynomial approximations for continuous functions over a closed interval. When used for smoothing an empirical CDF, Bernstein operators (see, e.g., Bustamante,, 2017) produce estimates that automatically respect the boundaries, mapping $[0,1]$ into $[0,1]$, and are genuine CDFs, since they are monotone non-decreasing and bounded between 0 and 1 by construction. They also enjoy shape-preserving properties, meaning the smoothed curve inherits the qualitative shape constraints of a CDF. The idea of leveraging Bernstein polynomials for nonparametric estimation dates back to Vitale, (1975), who studied a Bernstein estimator for density functions on a bounded interval. Babu et al. , (2002) and Babu & Chaubey, (2006) later demonstrated the utility of Bernstein polynomials for CDF estimation, applying them to smooth the empirical CDF and the associated normalized histogram under strongly mixing data. Subsequent work has analyzed the statistical properties of Bernstein estimators in detail. Leblanc, (2012a, b) derived higher-order bias and variance expansions for the Bernstein CDF estimator and showed that it has asymptotically negligible boundary bias and can be more efficient than the (unsmoothed) empirical CDF. There have also been several developments and extensions of Bernstein-type smoothers. For example, Jmaei et al. , (2017); Slaoui, (2022) develop and analyze a recursive approach to Bernstein CDF estimation, including a Robbins–Monro style estimator and moderate deviation theory. Erdoğan et al. , (2019) introduced an alternative CDF estimator using rational Bernstein polynomials, which generalizes the classical Bernstein approach to improve flexibility. For distributions supported on $0,\infty)$, Hanebeck & Klar, (2021) proposed using Szasz–Mirakyan positive linear operators, an analog of Bernstein operators for $0,\infty)$, to construct smooth CDF estimators for lifetime or survival data, and they showed that this method outperforms the empirical CDF in MSE and MISE as well. Moreover, Bernstein polynomials have been adapted to other complex settings. For example, Belalia et al. , (2017) estimated conditional CDFs by smoothing regression estimators, and Khardani, (2024) extended Bernstein-polynomial CDF and quantile estimation to right-censored data, highlighting their broad applicability. All of these approaches share a common goal of producing a smoother and more efficient estimate of the CDF while enforcing the shape constraints that are intrinsic to CDFs.
@@
Despite the rich literature on handling missing data and on smooth CDF estimation separately, relatively little attention has been given to smoothing CDF estimators in the presence of missing data. This work aims to fill that gap by developing and analyzing a nonparametric CDF estimator for MAR data that marries the IPW principle with Bernstein polynomial smoothing. In other words, the IPW-adjusted empirical CDF, which corrects bias due to MAR missingness, is smoothed using the Bernstein operator (see (2.1) below) to obtain a smooth and shape-constrained estimate. This approach inherits the double benefit of unbiasedness under MAR, from the weighting, and reduced variance plus automatic boundary adaptation, from the Bernstein smoothing.
@@
The remainder of the paper is organized as follows. Section 2 introduces the statistical framework, defines the Bernstein operator, formulates the MAR setting, describes propensity-score estimation from discrete covariates, and constructs the Bernstein-smoothed IPW estimators $\smash{\widetilde{F}_{n,m}}$ (when propensities are known) and $\smash{\widehat{F}_{n,m}}$ (when propensities are estimated). Section 3 presents the results: pointwise bias and variance, optimal choices of $m$ based on MSE and MISE, and asymptotic normality. Results are given first for the pseudo estimator with known propensities (Section 3.1) and then for the feasible estimator with estimated propensities (Section 3.2), where we quantify the variance reduction from estimating the propensity. Section 4 reports a Monte Carlo study comparing the smoothed and unsmoothed estimators across sample sizes, including the integrated-IPW KDE of Dubnicka, (2009) as a benchmark. Section 5 applies the feasible Bernstein-smoothed estimator to a fasting plasma glucose dataset. Section 6 summarizes the contributions and outlines directions for future research. All proofs are collected in Section 7. For reproducibility, A links to the R code used to generate the figures, the simulation results, and the real-data application, and B provides a list of acronyms used throughout.
@@ 2 The setup
Consider the following setting. In the population under study, there are a continuous response variable $Y$ subject to missingness and a discrete auxiliary variable $\boldsymbol{X}=(X_{1},\ldots,X_{d})$, for some $d\in\mathbb{N}$, which is fully observed. This is a common scenario in practice, for example with categorical demographics in surveys. Take a random sample of $n$ independent and identically distributed (iid) units from the population. For each unit $i\in\{1,\ldots,n\}$ in the sample, denote by $Y_{i}$ and $\boldsymbol{X}_{i}=(X_{i1},\ldots,X_{id})$ the $i$th observation of the response and auxiliary variables, respectively. Let the Bernoulli random variable
@@
$\pi_{i}=\pi_{i}(\boldsymbol{X}_{i})=\mathsf{P}\left(\delta_{i}=1\mid Y_{i},\boldsymbol{X}_{i}\right)=\mathsf{P}\left(\delta_{i}=1\mid\boldsymbol{X}_{i}\right).$
@@
The first version assumes propensities are known, which is the case, for example, when the missingness mechanism is designed. In this case, $F$ can be estimated by the following IPW pseudo-empirical estimator:
$\widetilde{F}_{n}(y)=n^{-1}\sum_{i=1}^{n}\frac{\delta_{i}}{\pi_{i}(\boldsymbol{X}_{i})}\mathds{1}_{\{Y_{i}\leq y\}}.$ (2.2)
The second and more realistic version addresses the fact that, in practice, the propensity scores are usually unknown. The discreteness of $\boldsymbol{X}$ simplifies propensity score estimation for MAR adjustment, allowing us to rely on the following nonparametric estimator:
@@
For continuous auxiliary data, this could be a Nadaraya–Watson estimator; see Dubnicka, (2009, Eq. (4)). The CDF $F$ can then be estimated using the feasible empirical estimator:
@@
$\widetilde{F}_{n,m}(y)$ $=\mathcal{B}_{m}(\widetilde{F}_{n})(y)=\sum_{k=0}^{m}\widetilde{F}_{n}(k/m)b_{m,k}(y)=n^{-1}\sum_{k=0}^{m}\sum_{i=1}^{n}\frac{\delta_{i}}{\pi_{i}(\boldsymbol{X}_{i})}\mathds{1}_{\{Y_{i}\leq k/m\}}b_{m,k}(y),$ (2.4)

### v2

@@ Abstract
We study nonparametric estimation of univariate cumulative distribution functions (CDFs) pertaining to data missing at random. The proposed estimators smooth the inverse probability weighted (IPW) empirical CDF with the Bernstein operator, yielding monotone, $[0,1]$-valued curves that automatically adapt to bounded supports. We analyze two versions: a pseudo estimator that uses known propensities and a feasible estimator that uses propensities estimated nonparametrically from discrete auxiliary variables, the latter scenario being much more common in practice. For both, we derive pointwise bias and variance expansions, establish the optimal polynomial degree $m$ with respect to the mean integrated squared error, and prove the asymptotic normality. A key finding is that the feasible estimator has a smaller variance than the pseudo estimator by an explicit nonnegative correction term. We also develop an efficient degree selection procedure via least-squares cross-validation. Monte Carlo experiments show that, for small to moderate sample sizes, the Bernstein-smoothed pseudo and feasible estimators outperform their unsmoothed counterparts and the integrated version of the IPW kernel density estimator proposed by Dubnicka (2009), under certain models. A real-data application to fasting plasma glucose from the 2017–2018 NHANES survey illustrates the method in a practical setting. All code needed to reproduce our analyses is readily accessible on GitHub.
@@ 1 Introduction
The cumulative distribution function (CDF) and its inverse (the quantile function) are fundamental objects in probability and statistics, providing a complete description of a random variable’s distribution. They offer insights into the entire distribution, including tail behavior, that summary measures like the mean or variance cannot capture. Accurate estimation of CDFs is crucial across numerous fields, for example, in survival analysis (where the CDF relates to survival probabilities), reliability engineering (where the CDF gives failure probabilities and reliability over time), economics (where the CDF characterizes income, wealth, or return distributions), and risk management (where the loss CDF underpins tail-risk metrics such as Value-at-Risk), because one often needs to understand the full distribution rather than just a single parameter.
@@
In practice, statistical analysis is frequently complicated by incomplete data. Missing observations are ubiquitous in large studies, clinical trials, and surveys. If data are missing in a nontrivial fashion, standard nonparametric estimators such as the empirical CDF computed from observed cases can be biased for the true distribution unless the data are missing completely at random, that is, the missingness mechanism is independent of the data values (Little and Rubin, 2019, p.14). A more plausible assumption in many contexts is missing at random (MAR), which posits that the probability that an observation is missing may depend on fully observed auxiliary variables but not on the value of the missing observation itself (Rubin, 1976). Under MAR, one can obtain unbiased estimates by appropriately reweighting or imputing the observed data. A prominent approach is inverse probability weighting (IPW), inspired by the Horvitz–Thompson estimator from survey sampling (Horvitz and Thompson, 1952). In IPW, each observed case is weighted by the inverse of its observation probability (called the propensity score), often estimated from covariates (Rosenbaum and Rubin, 1983), yielding a weighted empirical CDF (sometimes called a pseudo-empirical CDF). When propensity scores are known or consistently estimated, this IPW empirical CDF is asymptotically unbiased for the full-population CDF.
@@
Alternative nonparametric strategies for MAR data include imputation-based methods and hybrid approaches. For instance, Cheng and Chu (1996) proposed a kernel regression imputation to consistently estimate the CDF and quantiles, proving strong uniform consistency and asymptotic normality. Hybrid strategies, in particular augmented IPW estimators, have been developed to improve efficiency by combining weighting with outcome imputation. Wang and Qin (2010) introduced an augmented IPW empirical-likelihood approach for CDFs with missing responses that delivers asymptotic Gaussian limits for the process. These methods underscore that, under MAR assumptions, one can recover distributional information using either weighting or imputation, or both, without fully specifying outcome models. Furthermore, these estimation strategies have also been adapted to the more complex case of nonignorable missing data (NMAR). Related work by Ding and Tang (2018) considered IPW, regression imputation, and an augmented approach for distribution and quantile estimation under a semiparametric NMAR model, finding that all three can reach the same asymptotic variance under suitable conditions.
@@
However, the IPW-based empirical CDF, like the ordinary empirical CDF, is a step function. When the underlying distribution is continuous, it is often desirable to smooth such step-function estimates to improve efficiency, reduce mean squared error (MSE), and produce a smooth, continuous CDF estimate. Smoothing can yield substantial gains; some smooth CDF estimators have been shown to asymptotically outperform the raw empirical CDF in mean integrated squared error (MISE). Kernel smoothing is a popular approach in the density estimation context; see, e.g., Rosenblatt (1956); Parzen (1962), or Silverman (1986); Wand and Jones (1995) for book treatments. This approach extends naturally to CDFs by integrating the kernel density estimate, resulting in the classical smooth CDF estimator introduced by Nadaraya (1964), which replaces the indicator steps of the empirical CDF with integrated kernels. An alternative strategy involves direct regression smoothing of the empirical CDF. For example, Cheng and Peng (2002) proposed a local linear estimator that can achieve a smaller asymptotic MISE than the integrated kernel estimator, although it does not guarantee monotonicity. These kernel methods have been adapted to handle missing-data scenarios; for example, Dubnicka (2009) developed an IPW kernel density estimator (KDE) under MAR. For CDFs under MAR, a natural idea is to smooth the IPW variant using similar techniques. For instance, one of the competitors considered in our Monte Carlo study (Section 4) is an integrated version of Dubnicka’s estimator.
@@
Smoothing the CDF can improve pointwise variance and yield well-behaved estimates, but one must address the well-known boundary bias that traditional kernel estimators exhibit on bounded supports; see, e.g., Zhang et al. (2020). If the support of the distribution is [0,1], integrated-KDE CDF estimators tend to incur bias near 0 and 1 due to a spill-over effect created by the fixed kernel. To mitigate this, various boundary-correction techniques have been proposed, including reflecting the data at boundaries or using boundary kernels tailored for the edges (Tenreiro, 2013). A user-friendly and highly effective strategy employs asymmetric kernels, which inherently respect support limits and ensure monotonicity by construction; see, e.g., Lafaye de Micheaux and Ouimet (2021); Mansouri et al. (2023). Alternatively, monotonicity constraints can be imposed on smoothing techniques that do not naturally guarantee it. For instance, Xue and Wang (2010) proposed a constrained polynomial spline approach that smooths the empirical CDF while enforcing monotonicity, thereby guaranteeing a valid CDF estimate.
@@
In this paper, the focus is on CDFs supported on the unit interval $[0,1]$ (more general bounded supports can often be transformed to $[0,1]$). For such cases, Bernstein polynomials offer an elegant and effective smoothing technique. Bernstein polynomials (Bernstein, 1912) were originally introduced as a constructive proof of the Weierstrass approximation theorem (Weierstraß, 1885), which asserts the existence of uniform polynomial approximations for continuous functions over a closed interval. When used for smoothing an empirical CDF, Bernstein operators (see, e.g., Bustamante, 2017) produce estimates that automatically respect the boundaries, mapping $[0,1]$ into $[0,1]$, and are genuine CDFs, since they are monotone non-decreasing and bounded between 0 and 1 by construction. They also enjoy shape-preserving properties, meaning that the smoothed curve inherits the qualitative shape constraints of a CDF. The idea of leveraging Bernstein polynomials for nonparametric estimation dates back to Vitale (1975), who studied a Bernstein estimator for density functions on a bounded interval. Babu et al. (2002) and Babu and Chaubey (2006) later demonstrated the utility of Bernstein polynomials for CDF estimation, applying them to smooth the empirical CDF and the associated normalized histogram under strongly mixing data. Subsequent work has analyzed the statistical properties of Bernstein estimators in detail. Leblanc (2012a, b) derived higher-order bias and variance expansions for the Bernstein CDF estimator and showed that it has asymptotically negligible boundary bias and can be more efficient than the (unsmoothed) empirical CDF. There have also been several developments and extensions of Bernstein-type smoothers. For example, Jmaei et al. (2017) and Slaoui (2022) developed and analyzed a recursive approach to Bernstein CDF estimation, including a Robbins–Monro style estimator and moderate deviation theory. Erdoğan et al. (2019) introduced an alternative CDF estimator using rational Bernstein polynomials, which generalizes the classical Bernstein approach to improve flexibility. For distributions supported on $0,\infty)$, Hanebeck and Klar (2021) proposed using Szasz–Mirakyan positive linear operators, an analog of Bernstein operators for $0,\infty)$, to construct smooth CDF estimators for lifetime or survival data, and they showed that this method outperforms the empirical CDF in MSE and MISE as well. Moreover, Bernstein polynomials have been adapted to other complex settings. For example, Belalia et al. (2017) estimated conditional CDFs by smoothing regression estimators, and Khardani (2024) extended Bernstein-polynomial CDF and quantile estimation to right-censored data, highlighting their broad applicability. All of these approaches share a common goal of producing a smoother and more efficient estimate of the CDF while enforcing the shape constraints that are intrinsic to CDFs.
@@
Despite the rich literature on handling missing data and on smooth CDF estimation separately, relatively little attention has been given to smoothing CDF estimators in the presence of missing data. This work aims to fill that gap by developing and analyzing a nonparametric CDF estimator for MAR data that marries the IPW principle with Bernstein polynomial smoothing. In other words, the IPW-adjusted empirical CDF, which corrects bias due to MAR missingness, is smoothed using the Bernstein operator (see (2.1) below) to obtain a smooth and shape-constrained estimate. This approach inherits two benefits: unbiasedness under MAR from the weighting and reduced variance plus automatic boundary adaptation from the Bernstein smoothing.
@@
The remainder of the paper is organized as follows. Section 2 introduces the statistical framework, defines the Bernstein operator, formulates the MAR setting, describes propensity-score estimation from discrete covariates, and constructs the Bernstein-smoothed IPW estimators $\smash{\widetilde{F}_{n,m}}$ (when propensities are known) and $\smash{\widehat{F}_{n,m}}$ (when propensities are estimated). Section 3 presents the results: pointwise bias and variance, optimal choices of $m$ based on MSE and MISE, and asymptotic normality. Results are given first for the pseudo estimator with known propensities (Section 3.1) and then for the feasible estimator with estimated propensities (Section 3.2), where we quantify the variance reduction from estimating the propensity scores. Section 4 reports a Monte Carlo study comparing the smoothed and unsmoothed estimators across sample sizes, including the integrated-IPW KDE of Dubnicka (2009) as a benchmark. Section 5 applies the feasible Bernstein-smoothed estimator to a fasting plasma glucose dataset. Section 6 summarizes the contributions and outlines directions for future research. All proofs are collected in Section 7. For reproducibility, Appendix A links to the R code used to generate the figures, the simulation results, and the real-data application. Appendix B provides a list of acronyms used throughout.
@@ 2 The setup
Remark 1.
The Bernstein operator in (2.1) is defined for functions on $[0,1]$, so our theory is stated on that support. This is not restrictive for bounded responses: if a variable $Z$ is supported on $[a,b]$ with $a<b$, then the linearly rescaled variable $U=(Z-a)/(b-a)\in[0,1]$ can be analyzed by our method, and the original CDF is recovered by $F_{Z}(z)=F_{U}((z-a)/(b-a))$ for $z\in[a,b]$.
Consider the following setting. In the population under study, there are a continuous response variable $Y$ subject to missingness and a discrete auxiliary variable $\boldsymbol{X}=(X_{1},\ldots,X_{d})$, for some $d\in\mathbb{N}$, which is fully observed. This is a common scenario in practice, for example with categorical demographics in surveys. Take a random sample of $n$ independent and identically distributed (i.i.d.) pairs $(\boldsymbol{X}_{1},Y_{1}),(\boldsymbol{X}_{2},Y_{2}),\ldots,(\boldsymbol{X}_{n},Y_{n})$ from the population. For each unit $i\in\{1,\ldots,n\}$ in the sample, let $Y_{i}$ and $\boldsymbol{X}_{i}=(X_{i1},\ldots,X_{id})$ denote the $i$th response and $i$th covariate, respectively. Let the Bernoulli random variable
@@
$\pi(\boldsymbol{X}_{i})=\mathsf{P}\left(\delta_{i}=1\mid Y_{i},\boldsymbol{X}_{i}\right)=\mathsf{P}\left(\delta_{i}=1\mid\boldsymbol{X}_{i}\right).$
@@
The first version assumes that propensities are known, which is the case, for example, when the missingness mechanism is designed. In this case, $F$ can be estimated by the following IPW pseudo-empirical estimator:
$\widetilde{F}_{n}(y)=n^{-1}\sum_{i=1}^{n}\frac{\delta_{i}}{\pi(\boldsymbol{X}_{i})}\mathds{1}_{\{Y_{i}\leq y\}}.$ (2.2)
The second and more realistic version addresses the fact that, in practice, the propensity scores are usually unknown. In this paper, we restrict attention to auxiliary variables $\boldsymbol{X}$ with discrete finite support. This choice is mainly for expositional convenience: it avoids the tedious task of treating the discrete and continuous cases in parallel. In the discrete case, propensity score estimation for MAR adjustment is simplified, allowing us to rely on the following nonparametric estimator:
@@
If the auxiliary variable were continuous, one would simply replace (2.3) by a Nadaraya–Watson estimator; see Dubnicka (2009, Eq. (4)). The main asymptotic results stated in Section 3 would remain the same, but presenting both settings side by side would add substantial technical bookkeeping. For this reason, we formulate the theory only for the discrete case. The CDF $F$ can then be estimated using the feasible empirical estimator:
@@
$\widetilde{F}_{n,m}(y)$ $=\mathcal{B}_{m}(\widetilde{F}_{n})(y)=\sum_{k=0}^{m}\widetilde{F}_{n}(k/m)\,b_{m,k}(y)=n^{-1}\sum_{k=0}^{m}\sum_{i=1}^{n}\frac{\delta_{i}}{\pi(\boldsymbol{X}_{i})}\mathds{1}_{\{Y_{i}\leq k/m\}}b_{m,k}(y),$ (2.4)